Two cars are 456 miles/hour are traveling toward each other at the rates of 42miles/hour and 34 miles/hour in how many will they meet?
Answer by stanbon(75887)

Two cars are 456 miles/hour are traveling toward each other at the rates of 42miles/hour and 34 miles/hour in how many will they meet?
---------------------
Closing speed is 42+34 = 76 mph
---
Time to meet = 456/76 = 6 hrs.
==================================
Cheers,
Stan H.